# SpikeGS: Learning 3D Gaussian Fields from Continuous Spike Stream

### v1

@@ Abstract
A spike camera is a specialized high-speed visual sensor that offers advantages such as high temporal resolution and high dynamic range compared to conventional frame cameras. These features provide the camera with significant advantages in many computer vision tasks.
However, the tasks of 3D reconstruction and novel view synthesis based on spike cameras remain underdeveloped.
Although there are existing methods for learning neural radiance fields from spike stream, they either lack robustness in extremely noisy, low-quality lighting conditions or suffer from high computational complexity due to the deep fully connected neural networks and ray marching rendering strategies used in neural radiance fields, making it difficult to recover fine texture details. In contrast, the latest advancements in 3DGS have achieved high-quality real-time rendering by optimizing the point cloud representation into Gaussian ellipsoids. Building on this, we introduce SpikeGS, the first method to learn 3D Gaussian fields solely from spike stream. We designed a differentiable spike stream rendering framework based on 3DGS, incorporating noise embedding and spiking neurons. By leveraging the multi-view consistency of 3DGS and the tile-based multi-threaded parallel rendering mechanism, we achieved high-quality real-time rendering results. Additionally, we introduced a spike rendering loss function that generalizes under varying illumination conditions. Our method can reconstruct view synthesis results with fine texture details from a continuous spike stream captured by a moving spike camera, while demonstrating high robustness in extremely noisy low-light scenarios. Experimental results on both real and synthetic datasets demonstrate that our method surpasses existing approaches in terms of rendering quality and speed.
Keywords:
Spike camera 3D Gaussian splatting Novel View Synthesis 3D reconstruction
@@ 1 Introduction
Although some studies have explored the application of 3DGS and NeRF in a unique type of neuromorphic camera known as an event camera, characterized by differential sampling. Representative examples include Ev-NeRF [21], EventNeRF [44], and EvGGS [49], all of which introduce neural radiance fields or Gaussian fields derived specifically from event streams. However, the inherent lack of texture detail in event data limits the effectiveness of these methods, resulting in suboptimal outcomes. SpikeNeRF [63] and Spike-NeRF [14] have both explored methods for learning neural radiance fields from spike stream. However, the deep fully connected neural networks and ray marching-based rendering strategies used in neural radiance fields make it difficult to achieve high-quality real-time rendering. Additionally, Spike-NeRF [14] does not demonstrate robustness in extremely noisy, low-light scenarios.
@@
1) We proposed the first differentiable rendering framework that learns 3D Gaussian fields solely from spike stream (Fig. 3). SpikeGS (Fig. 1) exhibits high robustness in extremely noisy, low-quality lighting scenarios.
@@ 3 Method
The 3D Gaussian approach does not rely on neural radiance fields. instead, it represents the scene as a series of 3D Gaussian distributions. Based on the initialized sparse point cloud or randomly generated point cloud, a set of 3D Gaussians, defined as $\boldsymbol{G}$, is parameterized by its mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, 3D covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $\alpha\in\mathbb{R}$ and color $c\in\mathbb{R}^{3}$. $c$ is represented by spherical harmonics for view-dependent appearance. The distribution of each Gaussian is defined as:
@@ 3.2 Spike Signal Model
[…]
where $I{({x,y,t})}$ represents the light intensity at pixel $(x,y)$ at time $t$, and $\eta$ is the photoelectric conversion rate. We will directly use $I\left(t\right)$ to represent the luminance intensity to simplify our presentation. Due to the limitations of circuits, each spike is read out at discrete time $t$, $t\in T$($T=N$, $N$ indicates the window length). Thus, the output of the spike camera is a spatial-temporal binary stream S with $H\times W\times N$ size. Here, $H$ and $W$ are the height and width resolution of the sensor, respectively, and N is the temporal window size of the spike stream. In our experiments, we set N to 256. The process of spike emission can be represented by the following equation:
@@ 3.3 Spike Noise Model
Inspired by [64], we understand that due to significant circuit differences, the noise characteristics of spike cameras differ greatly from traditional cameras. The photodetectors in spike cameras receive photons from the scene, but even under uniform illumination, the photons striking the diodes are not constant. The difference between the number of photons at a given moment and the ideal number of photons is referred to as shot noise $N_{p}$. The dark current noise generated by the thermal diffusion of charge carriers and defects within or on the surface of a PN junction is denoted as $N_{d}$. Additionally, differences in photodiode characteristics and capacitance contribute to variations in pixel sensitivity to incident light intensity, resulting in photo-response non-uniformity noise $N_{rnu}$. Moreover, the temporal delay between the generation of the reset signal and the subsequent release of the spike signal introduces quantization noise $N_{q}$. If the temporal length of the spike stream is not long enough, truncation noise will appear in the process from the spike to the image. we define truncation noise as $N_{c}$
Therefore, based on the above, we can define the equation for pulse noise as:
@@ 3.4 SpikeGS
Establish the relationship between rendered pixel values and the real-world light intensity values $I$. Before establishing the relationship between rendering spike stream and real spike stream, we first establish their relationship in terms of light intensity. The goal is to estimate the intensity values corresponding to the real light of the scene. Unlike NeRF, which is a backward mapping process that operates on a per-pixel basis, emitting rays from 2D pixels and integrating along sampled points on the ray via volume rendering to synthesize the corresponding pixel values shown as Fig. 2. On the other hand, 3D Gaussians utilize forward mapping. In this forward mapping, 3D Gaussian ellipsoids are splatted onto the 2D image plane through a rasterization pipeline. Therefore, in models based on 3D Gaussians, we cannot directly establish the relationship between the pixel ray $r$ and the real-world light intensity values $I$. Therefore, we chose to establish the relationship between the view(pose) and the real-world light intensity values $I$. We denote the estimated intensity value corresponding to the real light of the scene as:
@@
$\left.L=\left({1-\lambda}\right)||\hat{S}(G,t)-S_{gt}{||}_{1}~{}+~{}\lambda\left({\hat{S}\left({G,t}\right),S_{gt}}\right)_{D-SSIM}\right.$ (13)
@@ 4.1 Experimental Settings
Benchmark Datasets. We used the spike dataset provided by SpikeNeRF[63] to evaluate our model. The synthetic dataset includes six scenes (chair, ficus, hotdog, lego, materials, and mic), with each scene comprising 100 images from different viewpoints. This synthetic dataset is generated based on NeRF’s synthetic dataset using the spike generator provided in [64]. The original size of the synthetic dataset is 800x800, which we resize to 400x400, generating a spike stream with a resolution of 400x400x256. The value 256 represents the temporal length of each spike stream. The real-world dataset includes four scenes (dolls, box, toys, grid), each with 35 images from different viewpoints. This dataset is recorded using a handheld spike camera, capable of capturing spike stream at a spatial resolution of 250x400 and a temporal resolution of 20 KHz.
Baselines. Due to the relative lack of methods for novel view synthesis based on spike camera, we compared our approach with four spike cameras image reconstruction methods: Spk2img+NeRF, Spk2img+GS, SpikeNeRF [63] and Spike-NeRF [14] (This paper is not open-source, therefore, we manually reproduced the method. Since the [14] paper does not include experiments on real-world datasets, we do not test [14] on real datasets for the sake of fairness). Spk2img (Spk2imgNet) [60] is a neural network-based learning method capable of recovering images from event streams. We combined Spk2imgNet with NeRF and 3D Gaussian models. Spk2imgNet first recovers multi-view images from the event stream, and then these images are input into the NeRF model and the 3D Gaussian model for novel view synthesis and comparison.
Training Details. For both synthetic and real datasets, the inputs consist of the corresponding spike stream and poses. Since 3D Gaussian inputs require an initial point cloud, we randomly generate the initial point cloud for the synthetic dataset. For the real dataset, we use the sparse point cloud (points3D) provided by SpikeNeRF and exported by COLMAP. Our experiments were conducted on a single NVIDIA 4090, with 20K iterations for Spike2img+GS and our model, and 200K iterations for Spk2img+NeRF and SpikeNeRF. Notably, our framework is approximately 15 times faster than SpikeNeRF, and its memory consumption is only half that of NeRF-based methods.
Experimental Evaluation Metrics. In the synthetic dataset, we use PSNR, SSIM [51], and LPIPS [59] as our experimental evaluation metrics to quantify the distance between the synthesized novel views and the ground truth RGB images. For the real dataset, which lacks corresponding ground truth images, we employ NIQE [15] [35] and BRISQUE [13] [34] as no-reference image quality evaluation metrics. Below, we present the quantitative and qualitative comparisons of our model on both synthetic and real datasets.
@@ 4.2 Quantitative and Qualitative Results
Table 1: Quantitative comparison on the real dataset. Each color shading indicates the best and second-best result. Due to the lack of ground truth in the real data, we use the no-reference quality assessment methods BRISQUE [13] [34] and NIQE [15] [35] to evaluate our results.
@@
Table 2: Quantitative comparison on the synthetic dataset. Each color shading indicates the best and second-best result, respectively.
@@
Qualitative results are demonstrated both on synthetic and real-world data. Fig. 4 presents the rendered images with different methods based on synthetic data (The visualized spike images are simulated by point clouds). Our method shows significant results with noisy spike inputs. The rendered images with real data are shown in Fig. 5. In the figure are four real scenes under different lighting conditions: "dolls" with high light intensity, "grid" and "toys" with moderate illumination, and "box" with the lowest illumination. Similarly, SpikeGS produces clearer images with fine textures than other methods.
@@ 4.3 Ablation Study
In this subsection, we perform ablation studies on each component of our framework. Specifically, we analyze the impact of our proposed 3D Gaussian-based spike generation pipeline and the spike rendering loss function. We denote the complete spike rendering loss function as $L_{s}$. $L_{I_{in}}$ represents directly using estimated light intensity and reconstructed images as supervision. $L_{1}$ denotes the supervision using only $L_{1}$ loss. We also investigate the effect of removing the noise embedding, represented by $D_{noise}$. $Full$ represents our complete model. Tables  3 (For the NIQE metric in $Full(L_{s})$, we select the data with N=128 to present the ablation experiments conducted on synthetic and real datasets, respectively. Overall, our full model outperforms all other configurations tested in both settings (Additional qualitative comparison of the ablation will be presented in the supplementary materials).
Table 3: Ablation Study on the synthetic dataset. Each color shading indicates the best and second-best result.
@@ 5 Conclusion
This paper introduces SpikeGS, the first work to learn 3D Gaussian fields solely from spike stream. we propose a novel rendering framework based on spike stream. We model the generation process of spike stream using 3DGS and embed the spike generation pipeline into the differentiable rasterization process of 3DGS, deriving the backpropagation accordingly. Additionally, we introduce a novel loss function for spike stream. Our model can recover clear novel views with fine details from extremely noisy spike stream under low-quality lighting conditions, using only the spike stream as supervision. We demonstrate the effectiveness of our approach on both synthetic and real datasets.

### v6

@@ Abstract
A spike camera is a specialized high-speed visual sensor that offers advantages such as high temporal resolution and high dynamic range compared to conventional frame cameras. These features provide the camera with significant advantages in many computer vision tasks. However, the tasks of novel view synthesis based on spike cameras remain underdeveloped. Although there are existing methods for learning neural radiance fields from spike stream, they either lack robustness in extremely noisy, low-quality lighting conditions or suffer from high computational complexity due to the deep fully connected neural networks and ray marching rendering strategies used in neural radiance fields, making it difficult to recover fine texture details. In contrast, the latest advancements in 3DGS have achieved high-quality real-time rendering by optimizing the point cloud representation into Gaussian ellipsoids. Building on this, we introduce SpikeGS, the method to learn 3D Gaussian fields solely from spike stream. We designed a differentiable spike stream rendering framework based on 3DGS, incorporating noise embedding and spiking neurons. By leveraging the multi-view consistency of 3DGS and the tile-based multi-threaded parallel rendering mechanism, we achieved high-quality real-time rendering results. Additionally, we introduced a spike rendering loss function that generalizes under varying illumination conditions. Our method can reconstruct view synthesis results with fine texture details from a continuous spike stream captured by a moving spike camera, while demonstrating high robustness in extremely noisy low-light scenarios. Experimental results on both real and synthetic datasets demonstrate that our method surpasses existing approaches in terms of rendering quality and speed.
@@ 1 Introduction
Although some studies have explored the application of 3DGS and NeRF in a unique type of neuromorphic camera known as an event camera, characterized by differential sampling. Representative examples include Ev-NeRF [21], EventNeRF [44], and EvGGS [49], all of which introduce neural radiance fields or Gaussian fields derived specifically from event streams. However, the inherent lack of texture detail in event data limits the effectiveness of these methods, resulting in suboptimal outcomes. The concurrent work [66] [67] explored spike stream reconstruction under high-speed motion cameras. SpikeNeRF [63] and Spike-NeRF [14] have both explored methods for learning neural radiance fields from spike stream. However, the deep fully connected neural networks and ray marching-based rendering strategies used in neural radiance fields make it difficult to achieve high-quality real-time rendering. Additionally, Spike-NeRF [14] does not demonstrate robustness in extremely noisy, low-light scenarios.
@@
1) We proposed a novel differentiable rendering framework that learns 3D Gaussian fields solely from spike stream (Fig. 3). SpikeGS (Fig. 1) exhibits high robustness in extremely noisy, low-quality lighting scenarios.
@@ 3 Method
The 3D Gaussian approach does not rely on neural radiance fields. instead, it represents the scene as a series of 3D Gaussian distributions. Based on the initialized sparse point cloud or randomly generated point cloud, a set of 3D Gaussians, defined as $\boldsymbol{G}$, is parameterized by its mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, 3D covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $\alpha\in\mathbb{R}$ and color $\in\mathbb{R}^{3}. cc$ is represented by spherical harmonics for view-dependent appearance. The distribution of each Gaussian is defined as:
@@ 3.2 Spike Signal Model
$A{({x,y,T})}={\int_{0}^{T}\eta}\cdot I{({x,y,t})}dt~mod~\Omega$ (7)
where I(x,y,t)I{({x,y,t})} represents the light intensity at pixel $(x,y)$ at time $t$, and $\eta$ is the photoelectric conversion rate. We will directly use $I\left(t\right)$ to represent the luminance intensity to simplify our presentation. Due to the limitations of circuits, each spike is read out at discrete time $t$, $t\in T$ ($T=Nt$, where t represents the unit time step and N is the size of the time window). Thus, the output of the spike camera is a spatial-temporal binary stream S with $H\times W\times N$ size. Here, $H$ and $W$ are the height and width resolution of the sensor, respectively, and N is the temporal window size of the spike stream. In our experiments, we set N to 256. The process of spike emission can be represented by the following equation:
@@ 3.3 Spike Noise Model
Inspired by [64], we understand that due to significant circuit differences, the noise characteristics of spike cameras differ greatly from traditional cameras. The photodetectors in spike cameras receive photons from the scene, but even under uniform illumination, the photons striking the diodes are not constant. The difference between the number of photons at a given moment and the ideal number of photons is referred to as shot noise $N_{p}$. The dark current noise generated by the thermal diffusion of charge carriers and defects within or on the surface of a PN junction is denoted as $N_{d}$. Additionally, differences in photodiode characteristics and capacitance contribute to variations in pixel sensitivity to incident light intensity, resulting in photo-response non-uniformity noise $N_{rnu}$. Moreover, the temporal delay between the generation of the reset signal and the subsequent release of the spike signal introduces quantization noise $N_{q}$. If the temporal length of the spike stream is not long enough, truncation noise will appear in the process from the spike to the image. we define truncation noise as $N_{c}$. Therefore, based on the above, we can define the equation for pulse noise as:
@@ 3.4 SpikeGS
Establish the relationship between rendered pixel values and the real-world light intensity values $I$. Before establishing the relationship between rendering spike stream and real spike stream, we first establish their relationship in terms of light intensity. The goal is to estimate the intensity values corresponding to the real light of the scene. Unlike NeRF, which is a backward mapping process that operates on a per-pixel basis, emitting rays from 2D pixels and integrating along sampled points on the ray via volume rendering to synthesize the corresponding pixel values shown as Fig. 2. On the other hand, 3D Gaussians utilize forward mapping. In this forward mapping, 3D Gaussian ellipsoids are splatted onto the 2D image plane through a rasterization pipeline (Each ellipsoid corresponds to multiple two-dimensional pixels). Therefore, in models based on 3D Gaussians, we cannot directly establish the relationship between the pixel ray $r$ and the real-world light intensity values $I$. Therefore, we chose to establish the relationship between the view(pose) and the real-world light intensity values $I$. We denote the estimated intensity value corresponding to the real light of the scene as:
@@
[…]
@@ 4.1 Experimental Settings
Benchmark Datasets. We used the spike dataset provided by SpikeNeRF[63] to evaluate our model. The synthetic dataset includes six scenes (chair, ficus, hotdog, lego, materials and mic), with each scene comprising 100 images from different viewpoints. This synthetic dataset is generated based on NeRF’s synthetic dataset using the spike generator provided in [64]. The original size of the synthetic dataset is 800x800. The real-world dataset includes four scenes (dolls, box, toys, grid), each with 35 images from different viewpoints. This dataset is recorded using a handheld spike camera, capable of capturing spike stream at a spatial resolution of 250x400 and a temporal resolution of 20 KHz.
Baselines. Due to the relative lack of methods for novel view synthesis based on spike camera, we compared our approach with four spike cameras image reconstruction methods: Spk2img+NeRF, Spk2img+GS, SpikeNeRF [63] and Spike-NeRF [14] (This paper is not open-source, therefore, we manually reproduced the method. Since the [14] paper does not include experiments on real-world datasets, we do not test [14] on real datasets for the sake of fairness). Spk2img (Spk2imgNet) [60] is a neural network-based learning method capable of recovering images from event streams. We combined Spk2imgNet with NeRF and 3DGS model. Spk2imgNet first recovers multi-view images from the event stream, and then these images are input into the NeRF model and the 3DGS model for novel view synthesis and comparison.
Training Details. For both synthetic and real datasets, the inputs consist of the corresponding spike stream and poses. Since 3D Gaussian inputs require an initial point cloud, we randomly generate the initial point cloud for the synthetic dataset. For the real dataset, we use the sparse point cloud (points3D) provided by SpikeNeRF and exported by COLMAP. Our experiments were conducted on a single NVIDIA 4090, with 20K iterations for Spike2img+GS and our model, and 200K iterations for Spk2img+NeRF, SpikeNeRF and Spike-NeRF. Notably, our framework is approximately 15 times faster than SpikeNeRF, and its memory consumption is only half that of NeRF-based methods.
Experimental Evaluation Metrics. In the synthetic dataset, we use PSNR, SSIM [51], and LPIPS [59] as our experimental evaluation metrics to quantify the distance between the synthesized novel views and the ground truth RGB images. For the real dataset, which lacks corresponding ground truth images, we employ NIQE [15] [35] and BRISQUE [13] [34] as no-reference image quality evaluation metrics. Below, we present the quantitative and qualitative comparisons of our model on both synthetic and real datasets. In the table, each color shading indicates the best and second-best result.
@@ 4.2 Quantitative and Qualitative Results
Table 1: Quantitative comparison for novel view synthesis on the real dataset.
@@
Table 2: Quantitative comparison for novel view synthesis on the synthetic dataset.
@@
Qualitative results are demonstrated both on synthetic and real-world data. Fig. 4 presents the rendered images with different methods based on synthetic data (The visualized spike images are simulated by point clouds). Our method shows significant results with noisy spike inputs. The rendered images with real data are shown in Fig. 5. In the figure are four real scenes under different lighting conditions. SpikeGS produces clearer images with fine textures than other methods.
@@ 4.3 Ablation Study
In this subsection, we perform ablation studies on each component of our framework. Specifically, we analyze the impact of our proposed 3D Gaussian-based spike generation pipeline and the spike rendering loss function. We denote the complete spike rendering loss function as $L_{s}$. $L_{I_{in}}$ represents directly using estimated light intensity and reconstructed images as supervision. $L_{1}$ denotes the supervision using only $L_{1}$ loss. We also investigate the effect of removing the noise embedding, represented by $D_{noise}$. $Full$ represents our complete model.
Table 3: Ablation Study for novel view synthesis on the synthetic dataset.
@@ 5 Conclusion
This paper introduces SpikeGS, the work to learn 3D Gaussian fields solely from spike stream. we propose a novel rendering framework based on spike stream. We model the generation process of spike stream using 3DGS and embed the spike generation pipeline into the differentiable rasterization process of 3DGS, deriving the backpropagation accordingly. Additionally, we introduce a novel loss function for spike stream. Our model can recover clear novel views with fine details from extremely noisy spike stream under low-quality lighting conditions, using only the spike stream as supervision. We demonstrate the effectiveness of our approach on both synthetic and real datasets.